OpenStudy (anonymous):

5v-4/10 = 4/5 PLZ HELP

OpenStudy (anonymous):

Is it (5v-4)/10=4/5 or 5v-(4/10)=4/5

OpenStudy (anonymous):

\[\frac{ 5v-4 }{ 10 } = \frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Still here just looking, I would probably multiple both sides by ten and get the variable off the fraction it's easy to work that way. \[5v-4=\frac{ 40 }{ 5 }\]

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

As you can see from there its just simplify the 40/5 to 8 then some light arithmetic

OpenStudy (anonymous):

so whats the answer i got 0.8

OpenStudy (anonymous):

Did you divide 4 by 5 or 40 by 5 remember we multiplied both sides by 10 so 4/5 x 10=40/5 which is 8